An exchange economy has two consumers, A & B, and two types of goods, x & y. A has an endowment
of x = 20 & y = 15, and B has an endowment x = 10 & y = 20. If after a trade A winds up with 18 units of
x and 22 units of y, then B must have (x,y) =
(a.) (10,20)
(b.) (12,20)
(c.) (12,13)
(d.) (10,13)

Graph and label it please, thanks!
Answer #1

Total endowment of X=20+10=30 and of Y=15+20=35

If A ends up with X=18 and Y=22 then B will be left with 30-18=12 X and Y=35-22=13

12,13

thus ans is C